Directions: Use the P-value Approach for all hypothesis tests. Follow all the steps used in hypothesis testing as delineated in the Guided Notes or the textbook and be sure to verify the conditions for each test. Include any StatCrunch, SurfStat, graphing calculator, etc.. output used in the process. I am assuming that all questions will be answered using technology-it is no necessary to include by-hand construction of the test statistics. Submitting test statistics only will not earn you full credit for that portion of the problem. Assume that all samples are randomly obtained statCrunch, Surfstat, graphing aechnology-it is notQuestion 4: A persons protein requirement is the amount of protein they must consume in a day to maintain a stable weight, muscle mass, etc. A group of researchers wanted to know if the mean protein requirement (in grams of protein per day per kilogram of body weight) is the same for young adult males and females. The protein requirement of 10 males and 11 females was determined. The protein requirements of the subjects are given below. (Data based orn summary statistics in CampbellWW, et. al., Dietary protein requirements of younger and older adults. American Journal of Clinical Nutrition. 2008.) Males:0.76 0.59 0.440.46 0.51 0.63 0.290.43 0.59 0.74 Females: 0.80 0.68 0.51 0.54 0.63 0.70 0.80 0.49 0.74 0.62 0.82 Is there a difference in the mean protein requirement of young adult males and females? Use α 0.025 and be sure to verify the requirements for the test
Answer #1

Solved using PHStat:

fA 1 Pooled-Variance t Test for the Difference Between Two Means 2 (assumes equal population variances) 4 Hypothesized Difference 5 Level of Significance Population 1 Males 7 Sample Size 8 Sample Mean 9 Sample Standard Deviation 10 11 Sample Size 12 Sample Mean 13 Sample Standard Deviation 14 15 Population 1 Sample Degrees of Freedom10-1 16 Population 2 Sample Degrees of Freedom11-1 17 Total Degrees of Freedom 18 Pooled Variance 19 Standard Error 20 Difference in Sample Means 21 t Test Statistic AVERAGE(Males) STDEV.S(Males) 1010 0.544 0.1465 Population 2 Females AVERAGE(Females) STDEV.S(Females) 0.6664 0.1188 Intermediate Calculations 9 10 19 0.0176 0.0580 0.1224 2.1113 = 9+10 ((C16* C9A2)+(C17*C13A2))/C18 -SORT(C19*(1/C7+1/C11)) = C8-C12 = (C21-C4)/C20 Two-Tail Test 23 24 Lower Critical Value 25 Upper Critical Value 26 p-Value 27 (TINV.2T(C5, C18)) T.INV.2T(C5, C18) T.DIST.2T(ABS(C22), C18) 2.4334 2.4334 0.0482 Do not reject the null hypothesis

As at alpha = 0.025 null hypothesis is not rejected. There is not enough evidence to conclude that there is a difference in the mean protein requirement of young adult males and females.